OpenStudy (anonymous):

Solve the following equations for x. Identify any extraneous solutions. Show all work.

OpenStudy (anonymous):

\[\sqrt{6x+5=\sqrt{53}}\]

OpenStudy (anonymous):

I am completely stuck on this can you olz help.

OpenStudy (anonymous):

plz*